# TabReD: A Benchmark of Tabular Machine Learning in-the-Wild

### v1

@@ 4.1 How do recent tabular DL advances transfer to a more realistic evaluation setting?
MLP 0.9500 0.6015 0.8545 0.2508 0.4820 0.5504 0.1622 1.5470 3.6 $\pm$ 1.6
SNN 0.9492 0.5996 0.8551 0.2858 0.4838 0.5544 0.1651 1.5649 5.5 $\pm$ 1.8
DCNv2 0.9392 0.5955 0.8466 0.2770 0.4842 0.5532 0.1672 1.5782 6.4 $\pm$ 2.1
ResNet 0.9469 0.5998 0.8493 0.2743 0.4825 0.5527 0.1625 1.5021 4.6 $\pm$ 1.7
FT-Transformer 0.9622 0.5775 0.8571 0.2440 0.4820 0.5542 0.1625 1.5104 3.5 $\pm$ 1.5
MLP-PLR 0.9621 0.5957 0.8568 0.2438 0.4812 0.5527 0.1616 1.5177 2.5 $\pm$ 1.3
TabR 0.9487 0.5943 0.8501 0.2820 0.4828 0.5514 0.1639 1.4666 4.8 $\pm$ 1.6
TabR (causal) 0.9522 0.5850 0.8484 0.2851 0.4825 0.5541 0.1637 1.4622 4.5 $\pm$ 1.5
Trompt 0.9546 0.5792 0.8381 0.2596 0.4834 0.5563 0.1652 1.5722 5.9 $\pm$ 1.5
OOD Robustness Methods
CORAL 0.9498 0.6004 0.8549 0.2645 0.4821 0.5498 0.1622 1.5591 3.9 $\pm$ 1.8
DFR 0.9499 0.6013 0.8545 0.2494 0.4819 0.5515 0.1626 1.5513 4.0 $\pm$ 1.7
GBDT and MLP with embeddings (MLP-PLR) are the overall best models on the TabReD benchmark. These findings suggest that numerical feature embeddings [20], which have shown success in academic datasets, maintain their utility in the new evaluation scenario.
FT-Transformer is a runner-up, however, it can be slower to train because of the attention module that causes quadratic scaling of computational complexity w.r.t. the number of features. The latter point is relevant for TabReD, since the TabReD datasets have more features than an average academic dataset. Thus, during hyperparameter tuning, trying 25 configurations of FT-Transformer takes two days on average on TabReD, compared to four hours for trying 100 configurations of MLP-PLR.
DCNv2, SNN, ResNet and Trompt are no better than the MLP baseline. Although Trompt showed promising results on a benchmark from Grinsztajn et al. [22], it failed to generalize to TabReD. Furthermore, efficiency-wise, Trompt is significantly slower than MLP, and even slower than FT-Transformer.
OOD Robustness. Both DFR and DeepCORAL do not improve upon the MLP baseline, in line with recent work by Gardner et al. [18], Kolesnikov [35] for other distribution shifts.
Retrieval-Based Models prove to be less performant on TabReD. One notable exception is the Weather dataset, where TabR demonstrated the best result. When describing TabR in section 4, we explained why TabReD can be a challenging benchmark for retrieval-based models, and outlined potential improvements for such models, but we leave them for future work.
Less Pronounced Performance Margins. Overall, we observe less pronounced performance differences from architecture improvements on TabReD. For example, the average percentage improvement of XGBoost over the MLP model is 3.1%, 2.7% and 1% on the Grinsztajn et al. [22], Gorishniy et al. [21] and TabReD benchmarks respectively. Tree-based methods have the smallest margin on TabReD, despite the fact that most data preparation pipelines of these datasets were tailored with GBDT models (a popular choice on Kaggle and in many production use cases). Similar relation holds for relative gains of the FT-Transformer and numerical feature embeddings.
4.2 Influence of Data Validation Splits on Model Ranking
In this section, we attempt to answer Q2. For this we create three time-based splits (with a sliding window over all samples) and three corresponding randomly shuffled splits, keeping train, validation and test sizes the same. We average all results over 15 random initialization seeds and three data splits. We are interested in methods ranking and relative performance differences between random and time-based test splits. For this experiment, we consider MLP, MLP-PLR, XGBoost and TabR-S. This selection covers multiple different paradigms: retrieval-based models, parametric DL and strong non-DL baselines, and diverse performance on our benchmark (see Table 2). Results are summarized in Figure 1. Below, we highlight key takeaways.
Temporal Shift Influence on Performance. We see that the spread of scores depending on random initialization for each model and data split is generally larger on the time-split based test set (most clear on Sberbank Housing, Delivery ETA and Ecom Offers). This indicates that temporal shift is present in the proposed datasets. As the random splits are commonplace in current academic benchmarks, they could present overly optimistic performance estimates to what one might expect in real-world application scenarios.
Temporal Shift Influence on Ranking and Relative Difference. We can see that the ranking of different model categories and the spreads between model performance scores change when we use randomly split test sets. One notable example is XGBoost decreasing its performance margin to MLPs when evaluated on temporally shifted test sets (Sberbank Housing, Cooking Time, Delivery ETA, Weather, Ecom Offers and Quote Conversion datasets – most clearly seen comparing MLP (PLR) with XGBoost). This might indicate that GBDTs are less robust to shifts, conversely performing better on random splits, by possibly exploiting time-based leakage. Another notable example is TabR-S outperforming the baseline MLP (Cooking Time and Homesite Insurance) and even XGBoost (Weather). These findings prove that curating datasets that represent real-world use cases is important for the continued stable progress in the field of tabular machine learning and further research adoption.
5 Limitations and Future Work
We recognize that the TabReD benchmark is biased towards industry-relevant ML applications, with large sample sizes, extensive feature engineering and temporal data drift. Also, TabReD does not cover some important domains such as medicine, science, and social data. Finally, we note that the lack of precise feature information may limit some potential future applications of these datasets, like leveraging feature names and descriptions with LLMs.
We provide initial experiments on TabReD demonstrating the importance of taking temporal shifts into account and benchmarking a wide range of tabular DL algorithms and even OOD robustness methods. There are still many research questions and techniques like continual learning, gradual temporal shift mitigation, missing data imputation and feature selection, that could be explored with TabReD.
Another important future research direction is creating a clean set of smaller/simpler academic datasets, taking the issues like data leakage and limiting the use of non-tabular datasets, which we discuss in subsection 3.1. As we demonstrated, some progress on academic benchmarks transfers to the in-the-wild settings present in TabReD, and simple academic benchmarks could still be useful for fast exploration.
6 Conclusion
In this work, we introduced the TabReD benchmark, a collection of eight industry-grade tabular datasets with temporal train/validation/test splits and feature engineering to better represent real-world tabular ML deployment scenarios. Our experiments show that accounting for temporal shifts during evaluation significantly impacts model performance and ranking. We evaluated recent tabular DL methods on TabReD datasets to find out how well they perform in the presence of time-based data splits and more features. The results highlight MLP with embeddings for continuous features [20] as a simple and effective DL baseline. More advanced DL models, however, demonstrated less convincing performance in the considered setting.
In summary, we demonstrate the need for representative benchmarks for enabling robust evaluation and progress in tabular ML. TabReD is an initial step toward realistic, task-relevant benchmarks. We hope our work steers the field toward tackling substantively important real-world problems using properly vetted protocols.
Broader Impact. This paper aims to contribute to the progression of Tabular Machine Learning research. A key focus of this work is the benchmarking of Tabular ML methods in academic settings. If successful, this research could potentially accelerate the transition of academic findings into real-world applications, which could have both positive and negative impacts on society. However, we firmly advocate for cautious and robust advancements in this field. We believe that such progress would be positive in the end.
Appendix A Benchmark Details
We provide the datasheet in the datasheet.md file in the accompanying source code. All datasets are available on Kaggle.
Appendix B Experimental Setup Extended
B.1 Source Code
We include source code for reproducing the results in the supplementary material archive with a brief README.md with instructions on reproducing the experiments. We publish code for dataset preparation for Kaggle-based datasets and publish new raw datasets on the Kaggle platform.
B.2 Tuning, Evaluation, Model comparisons
In tuning and evaluation setup, we closely follow the procedure described in [19, 21].
@@ B.2 Tuning, Evaluation, Model comparisons
When comparing the models we take standard deviations over 15 random initializations into account. We rank method A below method B if $|\text{B}_{\text{mean}}-\text{A}_{\text{mean}}|<\text{B}_{\text{stddev}}$ and B score is better. We run hyperparameter optimization for 100 iterations for most models, the exceptions are FT-Transformer (which is significantly less efficient on datasets with hundreds of features) where we were able to run 25
@@
For the exact hyperparameter ranges see the source code. Tuning configs (exp/**/tuning.toml) together with the code are always the main sources of truth.
B.3 Additional Implementation Details
@@
We have taken each method’s implementation from the respective official code sources, except for Trompt, which doesn’t have an official code repository yet and instead was reproduced by us according to the information in the paper. We use default hyperparameters for the Trompt model from the paper [12].
For the DFR [32] baseline we finetune the last layer on the last 20% of datapoints.
@@
For CORAL [48] we define domains by splitting instances based on a timestamp variable into 9 chunks.
Appendix C Kaggle Competitions Table
@@
We provide a annotated-kaggle-competitions.csv with annotated Kaggle competitions in the supplementary repository. We used this table during sourcing the datasets from Kaggle. There are minimal annotations and notes. We annotate only tabular competitions with more than 500 competitors. We provide annotations for non-tabular datasets in this table.
Appendix D TabReD Datasets Overview
In this section, we give a basic overview for each dataset from the proposed benchmark.
Homesite Insurance
Task: Classification
#Samples: 260,753     #Features: 299     Year: 2015
Comments: This is a dataset from a Kaggle competition hosted by Homesite Insurance [14]. The task is predicting whether a customer will buy a home insurance policy based on user and insurance policy features (user, policy, sales and geographic information). Each row in the dataset corresponds to a potential [customer, policy] pair, the target indicates whether a customer bought the policy. Timeframe from 2013-01-01 to 2015-05-18.
Ecom Offers
Task: Classification
#Samples: 160,057     #Features: 119     Year: 2014
Comments: This is a dataset from a Kaggle competition hosted by the online book and game retailer DMDave [15]. The task in this dataset is a representative example of modeling customer loyalty in e-commerce. Concretely, the task is classifying whether a customer will redeem a discount offer based on features from two months’ worth of transaction history. We base our feature engineering on one of the top solutions [40]. Timeframe from 2013-03-01 to 2013-04-30.
Homecredit Default
Task: Classification
#Samples: 1,526,659     #Samples Used 381,664     #Features: 696     Year: 2024
Comments: This is a second iteration of the popular HomeCredit tabular competition [23]. The task is to predict whether bank clients will default on a loan, based on bank internal information and external information like credit bureau and tax registry data. This year competition focus was the model prediction stability over time. Compared to the more popular prior competition, this time there is more data and the timestamps are available. We base feature engineering and preprocessing code on top solutions [31]. Timeframe from 2019-01-01 to 2020-10-05.
Sberbank Housing
Task: Regression
#Samples: 28,321     #Features: 392     Year: 2017
Comments: This dataset is from a Kaggle competition, hosted by Sberbank [3]. This dataset provides information about over 30000 transactions made in the Moscow housing market. The task is to predict the sale price of each property using the provided features describing each property condition, location, and neighborhood, as well as country economic indicators at the moment of the sale. We base our preprocessing code on discussions and solutions from the competition [4, 1]. Timeframe from 2011-08-20 to 2015-06-30.
Cooking Time
Task: Regression
#Samples: 12,799,642     #Samples Used 319,986     #Features: 192     Year: 2024
Comments: For this dataset, the task is to determine how long it will take for a restaurant to prepare an order placed in a food delivery app. Features are constructed based on the information about the order contents and historical information about cooking time for the restaurant and brand, the target is a logarithm of minutes it took to cook the placed order. Timeframe from 2023-11-15 to 2024-01-03.
Delivery ETA
Task: Regression
#Samples: 17,044,043     #Samples Used 416,451     #Features: 223     Year: 2024
Comments: For this dataset, the task is to determine the estimated time of arrival of an order from an online grocery store. Features are constructed based on the courier availability, navigation data and various aggregations of historical information for different time slices, the target is the logarithm of minutes it took to deliver an order. Timeframe from 2023-10-20 to 2024-01-25.
Maps Routing
Task: Regression
#Samples: 13,639,272     #Samples Used 340,981     #Features: 986     Year: 2024
Comments: For this dataset, the task is to predict the travel time in the car navigation system based on the current road conditions. The features are aggregations of the road graph statistics for the particular route and various road details (like speed limits). The target is the logarithm of seconds per kilometre. Timeframe from 2023-11-01 to 2023-12-04.
Weather
Task: Regression
#Samples: 16,951,828     #Samples Used 423,795     #Features: 103     Year: 2024
Comments: For this dataset, the task is weather temperature forecasting. This is a dataset similar to the one introduced in [38], except it is larger, and the time-based split is available and used by default. The features are from weather station measurements and weather forecast physical models. The target is the true temperature for the moment in time. Timeframe from 2022-07-01 to 2023-07-30.
@@ Appendix E Detailed Academic Datasets Overview
In this section we go through each dataset, and list its problems and prior uses in literature. We specify whether time-based splits should preferably be used for the dataset and whether it is available (e.g. datasets come with timestamps)

### v4

@@ C.2 Tuning, evaluation, and model comparisons
MLP 6 1 4 2 2 3 4 5
SNN 6 1 4 4 5 5 8 6
DCNv2 8 2 6 4 5 4 9 7
ResNet 7 1 6 4 3 4 5 3
FT-Transformer 1 3 4 1 2 5 4 3
MLP (PLR) 1 2 4 1 1 4 1 4
Trompt 4 3 6 4 4 5 8 6
MLP aug. 5 1 6 4 4 5 6 4
MLP aug. rec. 5 2 9 3 5 5 7 4
TabR 6 2 5 4 4 3 7 1
ModernNCA 5 3 5 4 3 3 5 3
CORAL 6 1 4 4 2 3 4 6
DFR 6 1 4 2 2 4 5 5
@@
We run hyperparameter optimization for 100 iterations for most models, the exceptions are FT-Transformer (which is significantly less efficient on datasets with hundreds of features) where we were able to run 25
@@
For the exact hyperparameter search spaces, please see the source code. Tuning configs (exp/**/tuning.toml) together with the code are always the main sources of truth.
C.3 Additional Implementation Details
@@
We have taken each method’s implementation from the respective official code sources, except for Trompt, which doesn’t have an official code repository yet and instead was reproduced by us according to the information in the paper. We use default hyperparameters for the Trompt model from the paper (Chen et al., 2023b).
For the DFR (Kirichenko et al., 2023) baseline, we finetune the last layer on the last 20% of datapoints.
@@
For CORAL (Sun & Saenko, 2016) we define domains by splitting instances based on a timestamp variable into 9 chunks.
Appendix D Kaggle Competitions Table
@@
We provide annotated-kaggle-competitions.csv with annotated Kaggle competitions in the supplementary repository. We used this table during sourcing the datasets from Kaggle. There are minimal annotations and notes. We annotate only tabular competitions with more than 500 competitors. We provide annotations for non-tabular datasets in this table.
@@ Appendix E Detailed Academic Datasets Overview
In this section, we go through each dataset, and list its problems and prior uses in literature. We specify whether time-based splits should preferably be used for the dataset and whether it is available (e.g. datasets come with timestamps).
We also adhere to the definition of tabular dataset proposed by Kohli et al. (2024) and annotate the existing datasets with one of five categories: raw data (the least tabular type), homogeneously extracted features HomE (e.g. similar features from one source, like image descriptors or features), heterogeneous extracted features HetE (different concepts, still extracted from one source) and semi-tabular/tabular (when there are multiple sources for HetE features).